OpenStudy (anonymous):

I really REALLY need someone to explain rate of consumption for me. None of the online walk throughs make any sense to me. The equation I was given was a generic one. 3A + 4B --> 5C I have to find the rate of consumption of B, and the rate of formation of C. Please explain!!

OpenStudy (kainui):

Just out of curiosity, what level of chemistry is this? For instance, is there any calculus like derivatives or integrals?

OpenStudy (anonymous):

I'm just about to do my final exam for grade twelve chemistry, if that helps.

OpenStudy (anonymous):

Calculus and integrals? I don't think so.

OpenStudy (kainui):

Alright, so do they give you any other information about the reaction here or is that all? From this all I can really say is that the rate at which the left hand side decreases is the same as the rate at which the right hand side increases. So for instance, let's say you have 3A and 4B right now sitting in your hand. They combine together and make exactly 5C. So you know that the rate at which the reactant goes away has to be the same as the rate that the product forms, so we divide out the amounts so in a way you can see that B must react slower than C because it's changing 4 mols of B in the same time it takes to change 5 mols of C. To do that in the same amount of time, that rate has to be faster, right? \[rate = \frac{-1}{3}[A]_{change}=\frac{-1}{4}[B]_{change}=\frac{1}{5}[C]_{change}\] So you can see that I've made the rates on the change of A and B negative since they're opposite of C's change. Remember, they're going away while we're forming C. If that doesn't clear it up, I'll help explain it better, unless there's more information they give you, I can't really say much more I think.

OpenStudy (anonymous):

Alright... Now why do they change into fractions? That's what confusing me.

OpenStudy (kainui):

Well remember this is the change of the concentration, not the concentration itself. So maybe I should use a different notation. \[\frac{-1}{4}\Delta B = \frac{1}{5}\Delta C=rate\] So now remember the concentration is what we can reasonably understand. 4 B's reacts at the same time as 5 C's. So if you think of it like distance = rate * time, then to go the same "distance" which is to say the same amount of reaction completed we need to multiply the rate of change of B by 4 amounts of B for ever 5 amounts of C. I'm sorry if that's not really making sense, I'm going to try to think about how I can explain it better. Perhaps just notice that if you take that above equation then that means if you multiply both sides of the equation by 4 and 5 you get: \[-5\Delta B = 4\Delta C\] See how 5 times the rate of decrease in B is 4 times the rate of increase of C? C must be faster than B because B is multiplied by a higher number but they're both equal.

OpenStudy (anonymous):

I had to read it out loud to myself to understand. I understand what you're trying to get across now. So, the answer to this question would be that equation? Or is there an actual number?

OpenStudy (kainui):

Well all I can really see here is that you can see the relative rate of reaction of one to the other, but without actually seeing anything more... You can't really say what the absolute reaction rate is. Yeah, I sort of confused myself a couple times typing that up, so I probably need to go review this myself lol.

OpenStudy (anonymous):

Well, it's just driving me insane cause I missed the week my class covered this, and my teacher never took the time to explain to me afterwards, even when I spoke to her. So now I'm scrambling, and online step by steps are so confusing. You've been a great help though, I'm still a bit foggy, but I think I'm on the right path